If you have 9.7 moles of O2 how many moles of water could you make?

C3H8+ 5O2 → 3CO2 + 4H2O


1
Expert's answer
2021-05-17T04:09:22-0400

Solution:

Balanced chemical equation:

C3H8 + 5O2 → 3CO2 + 4H2O

According to the equation above: n(O2)/5 = n(H2O)/4

Thus,

n(H2O) = 4 × n(O2) / 5 = (4 × 9.7 mol) / 5 = 7.76 mol

n(H2O) = 7.76 mol


Answer: 7.76 moles of water (H2O).